OpenStudy (anonymous):

Given that f(x) = 2x + 5 and g(x) = x − 7, solve for f(g(x)) when x = −3

jimthompson5910 (jim_thompson5910):

g(x) = x − 7 g(-3) = -3 − 7 ... replace x with -3 g(-3) = ??? ... tell me what you get

OpenStudy (anonymous):

-10

OpenStudy (jdoe0001):

\(\bf f(x) = 2x + 5 \qquad {\color{brown}{ g(x)}} = x - 7 \\ \quad \\ f(\quad {\color{brown}{ g(x)}}\quad )=2{\color{brown}{ g(x)}}+5\to 2({\color{brown}{ x-7}})+5 \\ \quad \\ x={\color{blue}{ 3}} \\ \quad \\ f(\quad {\color{brown}{ g({\color{blue}{ 3}})}}\quad )=2{\color{brown}{ g({\color{blue}{ 3}})}}+5\to 2({\color{brown}{ {\color{blue}{ 3}}-7}})+5\)

jimthompson5910 (jim_thompson5910):

now you plug that into f(x)

jimthompson5910 (jim_thompson5910):

f(x) = 2x + 5 f(-10) = 2*(-10) + 5 f(-10) = ???

OpenStudy (anonymous):

-15

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thanks :D

jimthompson5910 (jim_thompson5910):

no problem